OpenStudy (anonymous):

A record that regularly sells for $12.99 has a discount of $3.50. Using the formula, s = p − d, where s = sale price, p = regular price and d = discount, find the sale price of this record. $16.49 $15.99 $9.99 $9.49

OpenStudy (anonymous):

@Joel_the_boss

JoelTheBoss (joel_the_boss):

9.49

OpenStudy (cj49):

12.99-3.50=??

OpenStudy (anonymous):

9.49

OpenStudy (cj49):

thats the ans mate..

OpenStudy (anonymous):

Bill travels 250 miles in 25 hours on a bike trip from Boston to Cleveland. Using d = rt, what rate of speed does he travel? 20 mph 10 mph 25 mph 5 mph

JoelTheBoss (joel_the_boss):

i think it might be 10 miles per hour not 100 percent sure

OpenStudy (anonymous):

A record that regularly sells for $7.99 has a discount of $1.50. Using the formula, s = p − d, where s = sale price, p = regular price and d = discount, find the sale price of this record. $9.49 $10.49 $6.49 $5.99

OpenStudy (anonymous):

@cj49

OpenStudy (cj49):

7.99-150=?

OpenStudy (anonymous):

-142.01

OpenStudy (cj49):

sorry its 7.99-1.50

OpenStudy (anonymous):

6.49

OpenStudy (cj49):

thats the ans

OpenStudy (anonymous):

Automobiles have 12-volt systems. What is the current in a 12-volt circuit that has a resistance of 6 ohms? Use the formula i = v/r to find the resistance, where I is the current, V is the voltage, and R is the resistance. 72 amperes 2 amperes 4 amperes 6 amperes

OpenStudy (anonymous):

@cj49

OpenStudy (cj49):

i guess u have to find the voltage

OpenStudy (cj49):

wait its 12/6=?

OpenStudy (anonymous):

2

OpenStudy (cj49):

thats correct!!

OpenStudy (anonymous):

thanks

OpenStudy (cj49):

np mate...

OpenStudy (anonymous):

@Joel_the_boss give @cj49 a medel

OpenStudy (anonymous):

medal*

JoelTheBoss (joel_the_boss):

ok

OpenStudy (anonymous):

thanks for the help, I got 100%

JoelTheBoss (joel_the_boss):

lol your welcome